# From Scale to Speed: Adaptive Test-Time Scaling for Image Editing

### v1

@@ 2 Related Work
Test-time scaling in image generation enhances quality by extending inference time. Early work focuses on increasing denoising steps [TTS_add_steps_1, TTS_add_steps_2, TTS_add_steps_3, TTS_add_steps_efficiency_1]. Recently, inspired by Chain-of-Thought (CoT) in LLMs [NLP_CoT, NLP_CoT_ORM, NLP_CoT_PRM_1, NLP_CoT_PRM_2], Image-CoT has emerged as a promising paradigm. Noise scaling [TTS_Baseline, TTS_SANA] generates multiple samples by perturbing the noise and selects the best as the final result. However, its computational cost scales linearly. Subsequent work aims to improve this trade-off. Some methods enhance sample diversity through prompt-level intervention, such as rewriting [prompt_TTS_Yoonjin, T2IR1] or reflective updates [Reflection_DiT, GenRef_CoT_ICCV2025, Image_CoT]. Others adapt search algorithms [TTS_Baseline, classical_search_TTS, Tree_Sample_TTS, Video_TTS_EVO, TTS_search_ICML, TTS_greedy] (e.g., MCTS) to treat the reverse diffusion chain as a search trajectory and change the noise based on verifier scores. Recent methods [Video_TTS, Image_CoT, Video_TTS_beam_search, UniGen_TTS, ICEdit_arxiv2025] use MLLMs as verifiers to prune low-potential trajectories early. However, most Image-CoT methods target T2I generation and are suboptimal for editing. To this end, we propose ADE-CoT, an edit-specific test-time scaling method, to address issues of inefficient resource allocation, unreliable early-stage verification, and redundant edited results during Image-CoT.
@@ 3.1 Difficulty-aware Resource Allocation
where $N_{\mathrm{min}}$ and $N$ respectively denote the minimal and original budgets, $S_{\mathrm{max}}$ is the maximum possible score, and $\gamma$ is a hyperparameter to control sensitivity. This formulation ensures that for easy edits where $S\rightarrow S_{\mathrm{max}}$, the budget $N_{a}$ converges to $N_{\mathrm{min}}$. Conversely, for difficult edits where $S\rightarrow 0$, $N_{a}$ approaches the original budget $N$. As a result, it allocates more computation to difficult cases and saves resources on easy ones. More details are in Supp. Sec. B.1.
@@ 3.2 Edit-specific Verification in Early Pruning
Here, $\sigma_{t_{e}}$ is the noise scale and $\epsilon_{\theta}(x_{t_{e}},T_{t_{e}})$ is the predicted noise. $x_{0|{t_{e}}}$ is the predicted clean latent, which is then decoded into the preview image $I_{0|{t_{e}}}$. In Supp. Sec. B.2.1, we show that this preview reflects the correctness of the final output, which provides the basis for edit-specific verifiers.
@@
where a higher $S_{\text{cap}}$ indicates better semantic alignment to the instruction. Detailed prompts are in Supp. Sec. B.2.4.
@@
Filtering visually similar candidates. Goal-directed image editing often yields multiple redundant edited results during large-scale sampling. Notably, this redundancy has been apparent in the early preview images (see Supp. Sec. B.2.5). To remove similar images, we extract visual embeddings from each preview using DINOv2 [DINO_v2] and compute the pairwise similarity between them. If the similarity score between two preview images exceeds a threshold $\tau_{\text{sim}}$, we discard the one with the lower evaluation score. This step ensures that visually distinct and high-potential candidates are retained for the subsequent generation stage.
Sorting by evaluated score. Finally, the remaining candidates are sorted by $S$ in descending order. We empirically find that candidates with higher early scores tend to achieve higher final scores (see Supp. Sec. C.3). This enables the opportunistic stopping stage to terminate search earlier.
@@ 3.3 Depth-first Opportunistic Stopping
Retaining top results. Motivated by the stronger correlation between final image quality and preview scores at later denoising stages (see Supp. Sec. C.3), we use an additional check at a later timestep $t_{l}$, where $t_{e}<t_{l}<T$. Similar to the early pruning stage, we generate a preview image for each candidate and compute its unified score at timestep $t_{l}$. Instead of a fixed threshold, we apply an adaptive filter to retain candidates with scores comparable to the current highest score. It dynamically prunes samples that are unlikely to be optimal. More details are in Supp. Sec. B.3.1.
Instance-specific verifier. While general scores $S_{\text{gen}}$ are effective for coarse-grained rank, they often assign the same top scores to many candidates (see Fig. 2(c)), even when some contain editing errors. This makes the final selection unreliable. To address this, we introduce an instance-specific verifier for fine-grained assessment. We found that a two-stage inquiry effectively guides the MLLM to notice critical details (see Sec. 4.2). Given $I_{\text{src}}$ and $c$, it first generates a set of yes-no questions about the current edit via prompt $P_{q}$, covering aspects such as instruction adherence and aesthetics. Then, it answers them using prompt $P_{a}$ to produce an instance-specific score $S_{\text{spec}}$ by counting yes responses, where yes indicates a correct edit of one aspect. We combine $S_{\text{spec}}$ with the unified score to penalize error candidates. The search process stops after $N_{\text{high}}$ candidates are intent-aligned (i.e., receiving all yes answers). Finally, the candidate with the highest score is selected as the output. Detailed prompts are in Supp. Sec. B.3.2.
@@ 4 Experiments
Implementation details. Our ADE-CoT is evaluated on three SOTA, open-sourced image editing models: Step1X- Edit [Step1X_Edit_arxiv2025], FLUX.1 Kontext [FLUX_Kontext_arxiv2025], and BAGEL [Bagel_2025_arxiv]. We use their default denoising steps, i.e., $T=28,28,50$. In the default setting, the early step $t_{e}$ is set to $8,8,16$ and the retain step $t_{l}$ is set to $16,16,36$. We adopt Qwen-VL-MAX [qwen25vl] for MLLM queries and VIE-Score [VIE_Score] as the general score $S_{\text{gen}}$. If not explicitly stated, we set $N_{\text{min}}=1,\gamma=0.15,S_{\text{max}}=10,\lambda_{\text{reg}}=1,\lambda_{\text{cap}}=3,S_{\text{rj}}=5,\tau_{\text{sim}}=0.98,N_{\text{high}}=4$. We generate five instance-specific yes-no questions per edit. To ensure robustness, all results are averages of three runs. More details are in Supp. C.1.
@@ 4.3 Ablation Study
What is the optimal way to obtain early previews? We compare three methods in Tab. 3: (1) extra denoising steps (as in TTS-EF [ICEdit_arxiv2025]), (2) directly decoding noisy latents, and (3) our one-step preview. Additional denoising steps produce high-quality previews but significantly increase NFE. Decoding noisy latents provides clear previews only at later stages, leading to the highest NFE. In contrast, our one-step preview generates clear images in early steps with no extra iterations, achieving comparable G_O with the lowest NFE across all models. Detailed analysis is in Supp. Sec. C.3.
@@ Appendix E Extended Related Work
Test-time scaling in image generation. Recent years have seen rapid development in generative models [tang2024crs, tang2025aerogen, Wang_2024, wang2025jasmine, wang2025editor, Bagel_2025_arxiv, FLUX_Kontext_arxiv2025, Step1X_Edit_arxiv2025, qwen_image_arxiv2025, li2025ld, xu2025scalar, lan2025flux, chen2025s, mao2025omni]. Test-time scaling, as a training-free approach, aims to improve generation quality by extending the inference time. Early work on diffusion-based models investigates this by scaling the number of denoising steps [TTS_add_steps_1, TTS_add_steps_2, TTS_add_steps_3, TTS_add_steps_efficiency_1, TTS_add_steps_efficiency_2, TTS_add_steps_efficiency_3]. More recently, following the success of Chain-of-Thought (CoT) [NLP_CoT, NLP_CoT_ORM, NLP_CoT_PRM_1, NLP_CoT_PRM_2, NLP_TTS_SETS, NLP_TTS_MC, bai2025univg, NLP_TTS_MCTS, ProAPO, yuan2025autodrive, chen2025finger, li2025next, yuan2025video], Image-CoT has emerged as a promising paradigm. The standard Image-CoT method is noise scaling [TTS_Baseline], which generates multiple candidates by perturbing the noise and selects the best image as the final output. While effective, its computational cost scales linearly with the number of candidates. Subsequent work aims to generate higher-quality candidates within a fixed budget. Some methods enhance candidate diversity through prompt-level intervention, which enhances prompts by rewriting [prompt_TTS_Yoonjin, T2IR1] or reflective updates for iterative refinement [Reflection_DiT, GenRef_CoT_ICCV2025, Image_CoT]. Another line of work adapts search algorithms to the diffusion process. These methods treat the reverse diffusion chain as a search trajectory [TTS_Baseline, classical_search_TTS, Tree_Sample_TTS, Video_TTS_EVO, TTS_search_ICML]. They change the noise based on verifier scores to select promising sampling directions. Recent methods [Video_TTS, Image_CoT, wu2025imagerysearch, UniGen_TTS, Video_TTS_beam_search, ICEdit_arxiv2025, TTS_SANA] utilize MLLMs as a verifier to prune low-potential trajectories at early denoising stages. However, most Image-CoT methods focus on text-to-image generation. They are inefficient for image editing due to three key issues: (1) fixed sampling budgets waste computation on simple edits, (2) general MLLM scores cause misjudgement during early pruning, and (3) large-scale sampling produces redundant correct outputs. To address these challenges, we propose ADE-CoT, an edit-specific test-time scaling method that incorporates difficulty-aware resource allocation, edit-specific verification, and depth-first opportunistic stopping.

### v3

@@ 2 Related Work
Test-time scaling in image generation enhances quality by extending inference time. Inspired by Chain-of-Thought (CoT) in LLMs [NLP_CoT, NLP_CoT_ORM, NLP_CoT_PRM_1, NLP_CoT_PRM_2], Image-CoT has emerged as a promising paradigm. Noise scaling [TTS_Baseline, TTS_SANA] generates multiple samples by perturbing the noise and selects the best as the final result. However, its computational cost scales linearly. Subsequent work aims to improve this trade-off. Some methods enhance sample diversity through prompt-level intervention, such as rewriting [prompt_TTS_Yoonjin, T2IR1] or reflective updates [Reflection_DiT, GenRef_CoT_ICCV2025, Image_CoT]. Others adapt search algorithms [TTS_Baseline, classical_search_TTS, Tree_Sample_TTS, Video_TTS_EVO, TTS_search_ICML, TTS_greedy] (e.g., MCTS) to treat the reverse diffusion chain as a search trajectory and change the noise based on verifier scores. Recent methods [Video_TTS, Image_CoT, Video_TTS_beam_search, UniGen_TTS, ICEdit_arxiv2025] use MLLMs as verifiers to prune low-potential trajectories early. However, most Image-CoT methods target T2I generation and are suboptimal for editing. To this end, we propose ADE-CoT, an edit-specific test-time scaling method, to address issues of inefficient resource allocation, unreliable early-stage verification, and redundant edited results during Image-CoT.
@@ 3.1 Difficulty-aware Resource Allocation
where $N_{\mathrm{min}}$ and $N$ respectively denote the minimal and original budgets, $S_{\mathrm{max}}$ is the maximum score, and $\gamma$ is a hyperparameter to control sensitivity. This formulation ensures that for easy edits where $S\rightarrow S_{\mathrm{max}}$, $N_{a}$ converges to $N_{\mathrm{min}}$. Conversely, for difficult edits where $S\rightarrow 0$, $N_{a}$ approaches $N$. As a result, it allocates more computation to difficult cases and saves resources on easy ones.
@@ 3.2 Edit-specific Verification in Early Pruning
Here, $\sigma_{t_{e}}$ is the noise scale and $\epsilon_{\theta}(x_{t_{e}},T_{t_{e}})$ is the predicted noise. $x_{0|{t_{e}}}$ is the predicted clean latent, which is then decoded into the preview image $I_{0|{t_{e}}}$. In Supp. B.2.1, we show that this preview reflects the correctness of the final output, which provides the basis for edit-specific verifiers.
@@
where a higher $S_{\text{cap}}$ indicates better semantic alignment to the instruction. Detailed prompts are in Supp. B.2.4.
@@
Filtering visually similar candidates. Goal-directed image editing often yields multiple redundant edited results during large-scale sampling. Notably, this redundancy has been apparent in the early preview images (see Supp.B.2.5). To remove similar images, we extract visual embeddings from each preview using DINOv2 [DINO_v2] and compute the pairwise similarity between them. If the similarity score between two preview images exceeds a threshold $\tau_{\text{sim}}$, we discard the one with the lower evaluation score. This step ensures that visually distinct and high-potential candidates are retained for the subsequent generation stage.
Sorting by evaluated score. Finally, the remaining candidates are sorted by $S$ in descending order. We empirically find that candidates with higher early scores tend to achieve higher final scores (see Supp. C.3). This enables the opportunistic stopping stage to terminate search earlier.
@@ 3.3 Depth-first Opportunistic Stopping
Retaining top results. Motivated by the stronger correlation between final image quality and preview scores at later denoising stages (see Supp. C.3), we use an additional check at a later timestep $t_{l}$, where $t_{e}<t_{l}<T$. Similar to the early pruning stage, we generate a preview for each candidate and compute its unified score at timestep $t_{l}$. Instead of a fixed threshold, we apply an adaptive filter to retain candidates with scores comparable to the current highest score. This dynamically prunes suboptimal samples.
Instance-specific verifier. While general scores $S_{\text{gen}}$ are effective for coarse-grained rank, they often assign the same top scores to many candidates (see Fig. 2(c)), even when some contain editing errors. This makes the final selection unreliable. To address this, we introduce an instance-specific verifier for fine-grained assessment. We found that a two-stage inquiry effectively guides the MLLM to notice critical details (see Sec. 4.2). Given $I_{\text{src}}$ and $c$, it first generates a set of yes-no questions about the current edit via prompt $P_{q}$, covering aspects such as instruction adherence and aesthetics. Then, it answers them using prompt $P_{a}$ to produce an instance-specific score $S_{\text{spec}}$ by counting yes responses, where yes indicates a correct edit of one aspect. We combine $S_{\text{spec}}$ with the unified score to penalize error candidates. The search process stops after $N_{\text{high}}$ candidates are intent-aligned (i.e., receiving yes answers). Finally, the highest-scoring candidate is selected as the output.
@@ 4 Experiments
Implementation details. Our ADE-CoT is evaluated on three SOTA, open-sourced image editing models: Step1X- Edit [Step1X_Edit_arxiv2025], FLUX.1 Kontext [FLUX_Kontext_arxiv2025], and BAGEL [Bagel_2025_arxiv]. We use their default denoising steps, i.e., $T=28,28,50$. In the default setting, the early step $t_{e}$ is set to $8,8,16$ and the retain step $t_{l}$ is set to $16,16,36$. We adopt Qwen-VL-MAX [qwen25vl] for MLLM queries and VIE-Score [VIE_Score] as the general score $S_{\text{gen}}$. We generate five instance-specific yes-no questions per edit. To ensure robustness, all results are averages of three runs. More details are in Supp. C.1.
@@ 4.3 Ablation Study
What is the optimal way to obtain early previews? We compare three methods in Tab. 3: (1) extra denoising steps (as in TTS-EF [ICEdit_arxiv2025]), (2) directly decoding noisy latents, and (3) our one-step preview. Additional denoising steps produce high-quality previews but significantly increase NFE. Decoding noisy latents provides clear previews only at later stages, leading to the highest NFE. In contrast, our one-step preview generates clear images in early steps with no extra iterations, achieving comparable G_O with the lowest NFE across all models. Detailed analysis is in Supp. C.3.
@@ Appendix E Extended Related Work
Test-time scaling in image generation. Recent years have seen rapid development in generative models [tang2024crs, tang2025aerogen, Wang_2024, wang2025jasmine, wang2025editor, Bagel_2025_arxiv, FLUX_Kontext_arxiv2025, Step1X_Edit_arxiv2025, qwen_image_arxiv2025, li2025ld, xu2025scalar, wang2026geometry, jin2026semanticcontextmattersimproving, lan2025flux, chen2025s, mao2025omni, yu2025frequency]. Test-time scaling, as a training-free approach, aims to improve generation quality by extending the inference time. Early work on diffusion-based models investigates this by scaling the number of denoising steps [TTS_add_steps_1, TTS_add_steps_2, TTS_add_steps_3, TTS_add_steps_efficiency_1, TTS_add_steps_efficiency_2, TTS_add_steps_efficiency_3]. More recently, following the success of Chain-of-Thought (CoT) [NLP_CoT, NLP_CoT_ORM, NLP_CoT_PRM_1, NLP_CoT_PRM_2, NLP_TTS_SETS, NLP_TTS_MC, bai2025univg, NLP_TTS_MCTS, ProAPO, yuan2025autodrive, chen2025finger, li2025next, yuan2025image, yuan2025video], Image-CoT has emerged as a promising paradigm. The standard Image-CoT method is noise scaling [TTS_Baseline], which generates multiple candidates by perturbing the noise and selects the best image as the final output. While effective, its computational cost scales linearly with the number of candidates. Subsequent work aims to generate higher-quality candidates within a fixed budget. Some methods enhance candidate diversity through prompt-level intervention, which enhances prompts by rewriting [prompt_TTS_Yoonjin, T2IR1] or reflective updates for iterative refinement [Reflection_DiT, GenRef_CoT_ICCV2025, Image_CoT]. Another line of work adapts search algorithms to the diffusion process. These methods treat the reverse diffusion chain as a search trajectory [TTS_Baseline, classical_search_TTS, Tree_Sample_TTS, Video_TTS_EVO, TTS_search_ICML]. They change the noise based on verifier scores to select promising sampling directions. Recent methods [Video_TTS, Image_CoT, wu2025imagerysearch, UniGen_TTS, Video_TTS_beam_search, ICEdit_arxiv2025, TTS_SANA] utilize MLLMs as a verifier to prune low-potential trajectories at early denoising stages. However, most Image-CoT methods focus on text-to-image generation. They are inefficient for image editing due to three key issues: (1) fixed sampling budgets waste computation on simple edits, (2) general MLLM scores cause misjudgement during early pruning, and (3) large-scale sampling produces redundant correct outputs. To address these challenges, we propose ADE-CoT, an edit-specific test-time scaling method that incorporates difficulty-aware resource allocation, edit-specific verification, and depth-first opportunistic stopping.